OpenStudy (anonymous):

A triangler lot has a perimeter of 1900feet. one side is 200feet longer than the shortes side while the third side is 500feet longer than the stortest side and the lengths of all three sides.

OpenStudy (anonymous):

let shortest side = x ft then second side = (x + 200) ft then third side = (x + 500) ft ATQ, x + (x + 200) + (x + 500) = 1900 3x + 700 = 1900 3x = 1900 - 700 x = 1200/3 x = 400 so shortest side = x ft = 400 ft second side = (x + 200) ft = (400 + 200) ft = 600 ft third side = (x + 500) ft = (400 + 500) ft = 900 ft